Kelly has $5000 to deposit into two accounts. one account earn 4.5% simple interest, the other account earns 7.5% simple interest. After 1 year, the total interest earned in both accounts is $363. How much was invested in each account ?

Answer by mananth(16949)


Part I 4.50% per annum ------------- Amount invested =x
Part II 7.50% per annum ------------ Amount invested = y =
5000
Interest----- 363

Part I 4.50% per annum ---x
Part II 7.50% per annum ---y
Total investment
x + 1 y= 5000 -------------1
Interest on both investments
4.50% x + 7.50% y= 363
Multiply by 100
4.5 x + 7.5 y= 36300.00 --------2
Multiply (1) by -4.5
we get
-4.5 x -4.5 y= -22500.00
Add this to (2)
0 x 3 y= 13800
divide by 3
y = 4600
Part I 4.50% $ 400
Part II 7.50% $ 4600

CHECK
400 --------- 4.50% ------- 18.00
4600 ------------- 7.50% ------- 345.00
Total -------------------- 363.00

m.ananth@hotmail.ca